A 3.2 Hz continuous wave travels on a Slinky. If the wavelength is 0.49 m, what is the speed of waves on the Slinky (in m/s)?
1
Expert's answer
2019-04-29T09:44:00-0400

"\lambda = \frac{\upsilon}{\nu}"

"\lambda" - wavelength, m

"\upsilon" - speed of the wave, m/s

"\nu" - frequency of the wave, Hz

"\upsilon = \lambda \cdot \nu"

"\upsilon = 0.49 \cdot 3.2 = 1.568 \; m/s \approx 1.6 \; m/s"


Answer: 1.6 m/s.